# GeoDynamics: A Geometric State‑Space Neural Network for Understanding Brain Dynamics on Riemannian Manifolds

### v1

@@ Abstract
State‑space models (SSMs) have become a cornerstone for unraveling brain dynamics, revealing how latent neural states evolve over time and give rise to observed signals. By combining deep learning’s flexibility with SSMs’ principled dynamical structure, recent studies have achieved powerful fits to functional neuroimaging data. However, most approaches still view the brain as a set of loosely connected regions or impose oversimplified network priors, falling short of a truly holistic, self‐organized dynamical system perspective. Brain functional connectivity (FC) at each time point naturally forms a symmetric positive–definite (SPD) matrix, which lives on a curved Riemannian manifold rather than in Euclidean space. Capturing the trajectories of these SPD matrices is key to understanding how coordinated networks support cognition and behavior. To this end, we introduce GeoDynamics, a geometric state‑space neural network that tracks latent brain‐state trajectories directly on the high‑dimensional SPD manifold. GeoDynamics embeds each connectivity matrix into a manifold‑aware recurrent framework, learning smooth, geometry‑respecting transitions that reveal task‐driven state changes and early markers of Alzheimer’s, Parkinson’s, and autism. Beyond neuroscience, we validate GeoDynamics on human action recognition benchmarks (UTKinect, Florence, HDM05), demonstrating its scalability and robustness in modeling complex spatiotemporal dynamics across diverse domains.
@@ 1 Introduction
The widespread success of recurrent neural networks (RNNs, Fig. 1 (a)) [60], including long short-term memory (LSTM) [31] and gated recurrent units (GRU) [14], in sequential modeling tasks such as natural language processing (NLP), has inspired numerous efforts to apply these architectures for characterizing brain dynamics [48; 49]. Recently, state space models (SSMs) (as shown in Fig.1 (b, black solid box)) [27; 26] have emerged as a powerful tool for capturing a system’s behavior using hidden variables, or “states", marked as $s_{t}$ (i.e., $s(t)$), which effectively model temporal dependencies in sequential data with well-established theoretical properties. These models have gained significant attraction in fields like computer vision (CV) and NLP due to their ability to represent complex temporal patterns. A more inclusive literature survey can be found in the Appendix A.1.
@@
Relevant work of SSM on brain functional studies. While state-space models (SSMs) have achieved success in CV and NLP applications, their use in brain functional studies has primarily focused on event-related fMRI data [22; 35], limiting their applicability to resting-state fMRI (rs-fMRI). To address this, [65] combined an auto-encoder for learning BOLD signal relationships with a hidden Markov model (HMM) for state transitions, but the separate training of these components reduced efficiency and neglected brain network spatial structures. Similarly, [67] proposed a linear SSM using variational Bayesian methods to infer effective connectivities from EEG and fMRI data. However, this approach struggles to capture the complex dynamics between evolving functional connectivity and underlying cognitive or behavioral outcomes.
@@
We have applied our proposed method to two types of system dynamics: brain dynamics and action recognition [6; 28]. While brain dynamics is our primary focus, action recognition serves as a validation task to assess the method’s generalization performance across different domains. In the application of understanding brain dynamics, upon which we refer to as GeoDynamics, we have evaluated model performance on the large-scale human brain connectome (HBC) databases – one Human Connectome Project [80] and four disease-related resting-state fMRI data: (1) Alzheimer’s Disease Neuroimaging Initiative (ADNI) [53], (2) Open Access Series of Imaging Studies (OASIS) [47], (3) Parkinson’s Progression Markers Initiative (PPMI) [51], and (4) the Autism Brain Imaging Data Exchange (ABIDE). For action recognition, we use three classic human action recognition (HAR) datasets including the Florence 3D Actions dataset [62], the HDM05 database [54], and the UTKinect-Action3D (UTK) dataset [76]. Our GeoDynamics has achieved significant results across both brain dynamics and action recognition tasks, demonstrating its effectiveness and practicality. These applications on both neuroscience and computer vision highlight the scalability and robustness of GeoDynamics in understanding complex spatio-temporal dynamics across diverse systems.
@@ 2.1 Classical State Space Model
The (continuous-time) state space model is
@@
where $s(t)\!\in\!\mathbb{R}^{N}$ is the system state, $x(t)\!\in\!\mathbb{R}$ is the control input, and $y(t)\!\in\!\mathbb{R}^{M}$ the output (we follow the single-input/single-output convention unless otherwise stated). The matrices $A\!\in\!\mathbb{R}^{N\times N}$, $B\!\in\!\mathbb{R}^{N\times 1}$, $C\!\in\!\mathbb{R}^{M\times N}$, and $D\!\in\!\mathbb{R}^{M\times 1}$ (often D=0D\!=\!0) encode transition and observation, as sketched in Fig. 1(b). While Eq. (1) assumes Euclidean variables, many neuroscientific signals—notably functional connectivity matrices—are naturally manifold-valued (SPD), motivating a geometry-aware extension.
@@ 2.2 Riemannian Geometry on the SPD Manifold
Let $\mathbb{G}=\mathrm{GL}(N)$ be the general linear group acting on $\mathcal{M}$ via $g\mathbf{.}X:=gXg^{\top}$. Under the Stein metric, such actions are isometric; restricting to the orthogonal subgroup $\mathbb{O}(N)\!\subset\!\mathbb{G}$ preserves distances, $d(g\mathbf{.}X,g\mathbf{.}Y)=d(X,Y)$ for all $X,Y\!\in\!\mathcal{M}$ (proof in Appendix A.2). We call this an isometric translation on $\mathcal{M}$ and denote it by
@@
so the learned parameter is $Z$ while $H$ is guaranteed SPD; consequently $X^{(k)}$ remains SPD (proof in Appendix A.3).
@@ 3.1 State Update and Observation on the SPD Manifold
Let $\tau$ denote the length of the temporal receptive field (i.e., sliding window). We model the evolution of $S^{(k)}$ and $Y^{(k)}$ at discrete time step $k$ through two components: Temporal aggregation of past states and inputs using the weighted Fréchet mean (wFM) to ensure intrinsic averaging along geodesics rather than Euclidean linear combinations; Group action translation on the manifold to represent state transitions and input effects in an isometry-preserving manner.
@@
For classification or task decoding, we employ a softmax layer:
@@
$\mathcal{L}=-\sum_{k=1}^{K}\sum_{q=1}^{Q}o_{kq}\log P(Q=q\mid\hat{y}^{(k)}),$ (13)
where $o_{kq}$ is the ground truth label.
@@
where $\epsilon>0$ is a stability constant. The nonlinearity is introduced via the exponential map $\exp(\cdot)$ on the Riemannian algebra, which guarantees that the output remains SPD after each convolutional block. Frobenius normalization is applied to control eigenvalue growth and maintain numerical stability during training.
@@ 3.2 SPD-Preserving Attention
$\delta(\cdot)=\frac{\exp\bigl([X*\mathcal{K}]\bigr)}{\max\bigl(\exp([X*\mathcal{K}])\bigr)},\qquad[X]:=\mathrm{diag}(\theta,X),$ (16)
where $\mathrm{diag}(\theta,X)$ pads the borders of the SPD matrix with zeros of size $\theta-1$ and introduces a small positive diagonal component to ensure strict positive definiteness (see Appendix A.5). The exponential operation $\exp(\cdot)$ plays a role analogous to the sigmoid function, smoothly mapping the weighted response into the interval $[0,1]$ while retaining manifold consistency. The resulting $\delta(\cdot)$ acts as a soft mask that adaptively modulates local connectivity patterns. The attention-weighted representation is then computed as
@@
Disease-relevant pattern localization. Many neurological and psychiatric disorders, including neurodegenerative diseases such as Alzheimer’s, are characterized by localized but propagating disruptions in functional connectivity. These disruptions often manifest as: (1) spatially specific abnormal hubs or subnetworks (e.g., default mode, limbic regions), (2) temporal shifts in network activation or coupling, (3) altered covariance structure in FC trajectories reflecting pathological propagation. In conventional SSM or GNN models, such disease-related signals may be smoothed out by global averaging or fixed-weight message passing. In contrast, the SPA mechanism learns spatially and temporally adaptive weighting functions directly on the SPD manifold, enabling: Enhanced sensitivity to local network deviations: the multiplicative modulation amplifies abnormal signal components (e.g., hyper- or hypo-connectivity in key subnetworks) while attenuating background fluctuations. Preservation of intrinsic geometry: since both the attention and the convolution are SPD-preserving, pathological patterns expressed as topological changes (e.g., altered covariance structure) remain valid on the manifold, avoiding distortions induced by Euclidean projections. Temporal adaptivity: $\delta(\cdot)$ is computed at each time step, enabling the network to detect evolving abnormal FC trajectories associated with disease progression. Importantly, the learned attention weights $\delta(\cdot)$ are anatomically and temporally interpretable: high attention scores highlight regions or subnetworks where deviations from healthy connectivity trajectories are most pronounced. This makes SPA not only an effective inductive bias for improving predictive performance, but also a tool for identifying candidate disease biomarkers.
@@ 3.3 Summary of the GeoDynamics Framework
The proposed GeoDynamics integrates: Geometric state evolution: Input, system states, and outputs are represented as SPD matrices evolving on $\mathcal{M}$, ensuring geometric consistency. Manifold-aware operators: Temporal aggregation via weighted Fréchet means and state transitions via orthogonal group actions guarantee that system trajectories remain on the manifold. Stable discretization: Matrix exponential schemes provide stable integration and controllable temporal dynamics. Efficient convolutional formulation: A global kernel expansion enables parallel training and scalable implementation. Geometric attention: SPA enhances interpretability and sensitivity to task-relevant regions.
@@ 4.1 Experimental Setup
SPD matrices construction. For HBC dataset. Each fMRI scan has been processed into $N$ mean time courses of BOLD signals, each with $T$ time points (where $N$ represents the number of brain parcellations), we employ a sliding window technique to capture functional brain dynamics. Specifically, we construct a $N\times N$ correlation matrix at each time point $t$ ($t=1,\dots,T$) based on the BOLD signal within the sliding window, centered at time $t$. This results in a sequence of FC matrices encoding the functional dynamics for each scan, represented as $\mathcal{X}=\{X(t)\mid t=1,\dots,T\}\in\mathbb{R}^{T\times N\times N}$, in Fig. 2 (a). For HAR dataset. HAR datasets exhibit variability due to differences in action duration, complexity, the number of action classes, and the technology used for data capture. Therefore, we first apply a preprocessing step following [57] to obtain the SPD matrices. This step involves fixing the root joint at the hip center (red dashed circle in Fig. 2 (b)) and calculating the relative 3D positional differences for all other N1N-1 joints. For each timestamp $t=1,\dots,T$, we obtain a $3\times(N-1)$-dimensional column vector $p(t)$ representing the relative displacements of the joints. Then, we compute covariance matrices using the method proposed in [57] to yield the SPD matrices. After that, we apply a sliding window technique to capture the dynamics over time, resulting in a sequence of SPD matrices $\mathcal{X}=\{X(t)\mid t=1,\dots,T\}\in\mathbb{R}^{T\times(3(N-1))\times(3(N-1))}$, as illustrated in Fig. 2 (b).
@@ 4.2 Results on Human Brain Connectivity Datasets
Interpretation. Unlike task-based fMRI, resting-state fMRI measures spontaneous neural activity, reflecting intrinsic connectivity rather than stimulus-driven dynamics. This may explain why spatial models perform competitively in ND classification tasks. From a biological standpoint, neurodegeneration primarily manifests as gradual network disconnection rather than transient dysfunction [56; 12]. Functional disruption often precedes measurable cognitive decline by many years [71]. The results in Table 3 thus highlight the clinical promise of deep learning models for detecting early ND-related alterations in large-scale brain networks.
Neuropsychiatric disorders. We further analyze Autism Spectrum Disorder (ASD) using the ABIDE dataset. As shown in Fig. 3 (blue colors), spatial models slightly outperform sequential ones, with SPDNet consistently ranking second only to our GeoDynamics. Both SPDNet and GeoDynamics share two key design principles: (1) manifold-based representation learning that preserves the intrinsic geometry of functional connectivity matrices (Fig. 2 (c)), and (2) spatio-temporal modeling that captures the evolution of connectivity patterns over time. The consistent advantage of these two approaches suggests that robust spatio-temporal representation, grounded in Riemannian geometry, is crucial for reliable diagnosis of neuropsychiatric conditions.
@@
Interpretation. Neuropsychiatric disorders such as Autism and Schizophrenia are characterized by atypical neural connectivity and abnormal variability in BOLD dynamics [59; 52; 39]. These alterations affect both spatial topology and temporal synchronization, suggesting that effective diagnostic models must jointly capture both aspects. In contrast, neurodegenerative diseases primarily disrupt large-scale connectivity due to neuronal loss, making static spatial features more informative. Integrating such disease-specific pathophysiological insights is therefore essential for model design and interpretation.

### v2

@@ Abstract
State‑space models (SSMs) have become a cornerstone for unraveling brain dynamics, capturing how latent neural states evolve over time and give rise to observed signals. By combining deep learning’s flexibility with SSMs’ principled dynamical structure, recent studies have achieved powerful fits to functional neuroimaging data. However, most approaches still view the brain as a set of loosely connected regions or impose oversimplified network priors, falling short of a truly holistic, self‐organized dynamical system perspective. Brain functional connectivity (FC) at each time point naturally forms a symmetric positive definite (SPD) matrix, which lives on a curved Riemannian manifold rather than in Euclidean space. Capturing the trajectories of these SPD matrices is key to understanding how coordinated networks support cognition and behavior. To this end, we introduce GeoDynamics, a geometric state space neural network that tracks latent brain state trajectories directly on the high‑dimensional SPD manifold. GeoDynamics embeds each connectivity matrix into a manifold‑aware recurrent framework, learning smooth, geometry‑respecting transitions that reveal task‐driven state changes and early markers of Alzheimer’s, Parkinson’s, and autism. Beyond neuroscience, we validate GeoDynamics on human action recognition benchmarks (UTKinect, Florence, HDM05), demonstrating its scalability and robustness in modeling complex spatiotemporal dynamics across diverse domains.
@@ 1 Introduction
The widespread success of recurrent neural networks (RNNs, Fig. 1 (a)) [59], including long short-term memory (LSTM) [31] and gated recurrent units (GRU) [14], in sequential modeling tasks such as natural language processing (NLP), has inspired numerous efforts to apply these architectures for characterizing brain dynamics [47; 48]. Recently, state space models (SSMs) (as shown in Fig.1 (b, black solid box)) [27; 26] have emerged as a powerful tool for capturing a system’s behavior using hidden variables, or “states", marked as $s_{t}$ (i.e., $s(t)$), which effectively model temporal dependencies in sequential data with well-established theoretical properties. These models have gained significant attraction in fields like computer vision (CV) and natural language processing (NLP) due to their ability to represent complex temporal patterns. A more inclusive literature survey can be found in the Appendix A.1.
@@
Relevant work of SSM on brain functional studies. While SSMs have achieved success in CV and NLP applications, their use in brain functional studies has primarily focused on event-related fMRI data [22; 35], limiting their applicability to resting-state fMRI (rs-fMRI). To address this, [64] combined an auto-encoder for learning BOLD signal relationships with a hidden Markov model (HMM) for state transitions, but the separate training of these components reduced efficiency and neglected brain network spatial structures. Similarly, [66] proposed a linear SSM using variational Bayesian methods to infer effective connectivities from EEG and fMRI data. However, this approach struggles to capture the complex dynamics between evolving functional connectivity and underlying cognitive or behavioral outcomes.
@@
We apply our proposed method (GeoDynamics) to two types of system dynamics: brain dynamics and action recognition [6; 28]. While brain dynamics is our primary focus, action recognition serves as a validation task to assess the method’s generalization performance across different domains. In the application of understanding brain dynamics, upon which we refer to as GeoDynamics, we have evaluated model performance on the large-scale human brain connectome (HBC) databases, including one Human Connectome Project [79] and four disease-related resting-state fMRI data: (1) Alzheimer’s Disease Neuroimaging Initiative (ADNI) [52], (2) Open Access Series of Imaging Studies (OASIS) [46], (3) Parkinson’s Progression Markers Initiative (PPMI) [50], and (4) the Autism Brain Imaging Data Exchange (ABIDE). For action recognition, we use three classic human action recognition (HAR) datasets including the Florence 3D Actions dataset [61], the HDM05 database [53], and the UTKinect-Action3D (UTK) dataset [75]. Our GeoDynamics has achieved significant results across both brain dynamics and action recognition tasks, demonstrating its effectiveness and practicality. These applications on both neuroscience and computer vision highlight the scalability and robustness of GeoDynamics in understanding complex spatio-temporal dynamics across diverse systems.
@@ 2.1 Classical State Space Model
The continuous-time state space model is defined as
@@
where $s(t)\!\in\!\mathbb{R}^{N}$ is the system state, $x(t)\!\in\!\mathbb{R}$ is the control input, and $y(t)\!\in\!\mathbb{R}^{M}$ the output (we follow the single-input/single-output convention unless otherwise stated). The matrices $A\!\in\!\mathbb{R}^{N\times N}$, $B\!\in\!\mathbb{R}^{N\times 1}$, $C\!\in\!\mathbb{R}^{M\times N}$, and $D\!\in\!\mathbb{R}^{M\times 1}$ (often D=0D\!=\!0) encode transition and observation, as sketched in Fig. 1(b). While Eq. (1) assumes Euclidean variables, many neuroscientific signals, notably functional connectivity matrices, are naturally manifold-valued (SPD), motivating a geometry-aware extension.
@@ 2.2 Riemannian Geometry on the SPD Manifold
Let $\mathbb{G}=\mathrm{GL}(N)$ be the general linear group acting on $\mathcal{M}$ via $g\mathbf{.}X:=gXg^{\top}$. Under the Stein metric, such actions are isometric, restricting to the orthogonal subgroup $\mathbb{O}(N)\!\subset\!\mathbb{G}$ preserves distances, $d(g\mathbf{.}X,g\mathbf{.}Y)=d(X,Y)$ for all $X,Y\!\in\!\mathcal{M}$ (proof in Appendix A.2). We call this an isometric translation on $\mathcal{M}$ and denote it by
@@
so the learned parameter is $Z$ while $H$ is guaranteed SPD, consequently $X^{(k)}$ remains SPD (proof in Appendix A.3).
@@ 3.1 State Update and Observation on the SPD Manifold
Let $\tau$ denote the length of the temporal receptive field (i.e., sliding window). We model the evolution of $S^{(k)}$ and $Y^{(k)}$ at discrete time step $k\in\{1,K\}$ through two components: (1) Temporal aggregation of past states and inputs using the weighted Fréchet mean (wFM) to ensure intrinsic averaging along geodesics rather than Euclidean linear combinations. (2) Group action translation on the manifold to represent state transitions and input effects in an isometry-preserving manner.
@@
For classification task and brain state decoding, we employ a softmax layer:
@@
$\mathcal{L}=-\frac{1}{E}\sum_{i=1}^{E}\sum_{q=1}^{Q}o_{iq}\log P(Q=q\mid\hat{y}_{i}^{(K)})$ (13)
where $o_{iq}$ is the ground truth label. $E$ denotes the number of subjects/samples.
@@
where $\epsilon>0$ is a stability constant. The nonlinearity is introduced via the exponential map $\exp(\cdot)$ on the Riemannian algebra, which guarantees that the output remains SPD after each convolutional block [15]. Frobenius normalization is applied to control eigenvalue growth and maintain numerical stability during training.
@@ 3.2 SPD-Preserving Attention
$\delta(\cdot)=\frac{\exp\bigl([X*\mathcal{K}]\bigr)}{\max\bigl(\exp([X*\mathcal{K}])\bigr)},\qquad[X]:=\mathrm{diag}(\rho,X),$ (16)
where $\mathrm{diag}(\rho,X)$ pads the borders of the SPD matrix with zeros of size $\rho-1$ and introduces a small positive diagonal component to ensure strict positive definiteness (see Appendix A.5). The exponential operation $\exp(\cdot)$ plays a role analogous to the sigmoid function, smoothly mapping the weighted response into the interval $[0,1]$ while retaining manifold consistency. The resulting $\delta(\cdot)$ acts as a soft mask that adaptively modulates local connectivity patterns. The attention-weighted representation is then computed as
@@
Disease-relevant pattern localization. Many neurological and psychiatric disorders, including neurodegenerative diseases such as Alzheimer’s, are characterized by localized but propagating disruptions in functional connectivity. These disruptions often manifest as: (1) spatially specific abnormal hubs or subnetworks (e.g., default mode, limbic regions), (2) temporal shifts in network activation or coupling, (3) altered covariance structure in FC trajectories reflecting pathological propagation. In conventional SSM or GNN models, such disease-related signals may be smoothed out by global averaging or fixed-weight message passing. In contrast, the SPA mechanism learns spatially and temporally adaptive weighting functions directly on the SPD manifold, enabling: Enhanced sensitivity to local network deviations: the multiplicative modulation amplifies abnormal signal components (e.g., hyper- or hypo-connectivity in key subnetworks) while attenuating background fluctuations. Preservation of intrinsic geometry: since both the attention and the convolution are SPD-preserving, pathological patterns expressed as topological changes (e.g., altered covariance structure) remain valid on the manifold, avoiding distortions induced by Euclidean projections. Temporal adaptivity: $\delta(\cdot)$ is computed at each time step, enabling the network to detect evolving abnormal FC trajectories associated with disease progression. Importantly, the learned attention weights $\delta(\cdot)$ are anatomically and temporally interpretable, i.e., high attention scores highlight regions or subnetworks where deviations from healthy connectivity trajectories are most pronounced. This makes SPA not only an effective inductive bias for improving predictive performance, but also a tool for identifying candidate disease biomarkers.
@@ 3.3 Summary of the GeoDynamics Framework
The proposed GeoDynamics integrates: (1) Geometric state evolution: Input, system states, and outputs are represented as SPD matrices evolving on $\mathcal{M}$, ensuring geometric consistency. (2) Manifold-aware operators: Temporal aggregation via weighted Fréchet means and state transitions via orthogonal group actions guarantee that system trajectories remain on the manifold. (3) Stable discretization: Matrix exponential schemes provide stable integration and controllable temporal dynamics. (4) Efficient convolutional formulation: A global kernel expansion enables parallel training and scalable implementation. (5) Geometric attention: SPA enhances interpretability and sensitivity to task-relevant regions.
@@ 4.1 Experimental Setup
SPD matrices construction. For HBC dataset. Each fMRI scan has been processed into $N$ mean time courses of BOLD signals, each with $T$ time points (where $N$ represents the number of brain parcellations), we employ a sliding window technique to capture functional brain dynamics. Specifically, we construct a $N\times N$ correlation matrix at each time point $t$ ($t=1,\dots,T$) based on the BOLD signal within the sliding window, centered at time $t$. This results in a sequence of FC matrices encoding the functional dynamics for each scan, represented as $\mathcal{X}=\{X(t)\mid t=1,\dots,T\}\in\mathbb{R}^{T\times N\times N}$, in Fig. 2 (a). For HAR dataset. HAR datasets exhibit variability due to differences in action duration, complexity, the number of action classes, and the technology used for data capture. Therefore, we first apply a preprocessing step following [56] to obtain the SPD matrices. This step involves fixing the root joint at the hip center (red dashed circle in Fig. 2 (b)) and calculating the relative 3D positional differences for all other N1N-1 joints. For each timestamp $t=1,\dots,T$, we obtain a $3\times(N-1)$-dimensional column vector representing the relative displacements of the joints. Then, we compute covariance matrices using the method proposed in [56] to yield the SPD matrices. After that, we apply a sliding window technique to capture the dynamics over time, resulting in a sequence of SPD matrices $\mathcal{X}=\{X(t)\mid t=1,\dots,T\}\in\mathbb{R}^{T\times(3(N-1))\times(3(N-1))}$, as illustrated in Fig. 2 (b).
@@ 4.2 Results on Human Brain Connectivity Datasets
Interpretation. Unlike task-based fMRI, resting-state fMRI measures spontaneous neural activity, reflecting intrinsic connectivity rather than stimulus-driven dynamics. This may explain why spatial models perform competitively in ND classification tasks. From a biological standpoint, neurodegeneration primarily manifests as gradual network disconnection rather than transient dysfunction [55; 12]. Functional disruption often precedes measurable cognitive decline by many years [70]. The results in Fig. 3 thus highlight the clinical promise of deep learning models for detecting early ND-related alterations in large-scale brain networks.
Neuropsychiatric disorders. We further analyze Autism Spectrum Disorder (ASD) using the ABIDE dataset. As shown in Fig. 3 (blue color), spatial models slightly outperform sequential ones, with SPDNet consistently ranking second only to our GeoDynamics. Both SPDNet and GeoDynamics share two key design principles: (1) manifold-based representation learning that preserves the intrinsic geometry of functional connectivity matrices (Fig. 2 (c)), and (2) spatio-temporal modeling that captures the evolution of connectivity patterns over time. The consistent advantage of these two approaches suggests that robust spatio-temporal representation, grounded in Riemannian geometry, is crucial for reliable diagnosis of neuropsychiatric conditions.
@@
Interpretation. Neuropsychiatric disorders such as Autism and Schizophrenia are characterized by atypical neural connectivity and abnormal variability in BOLD dynamics [58; 51; 38]. These alterations affect both spatial topology and temporal synchronization, suggesting that effective diagnostic models must jointly capture both aspects. In contrast, neurodegenerative diseases primarily disrupt large-scale connectivity due to neuronal loss, making static spatial features more informative. Integrating such disease-specific pathophysiological insights is therefore essential for model design and interpretation.